a rough road is banked at 30 degree with horizontal. the ratio of maximum to minimum safe speed for travelling on the road ia 

Grade:12th pass

1 Answers

Sushant Kumar
5 years ago
according to me value of coefficent of friction and radius is needed in the question
let coeffiecent of friction be \mu
Vmax=\sqrt{\frac{Rg(\tan\Theta +\mu )}{1-\mu \tan\Theta }}  where R=radius  g=accleration due to gravity(=9.8)  \Theta=30 degress
 
 
Vmin=\sqrt{\frac{Rg(\tan\Theta -\mu )}{1+\mu \tan\Theta }}
 
 
 
hence  you can find the ratio 
derivation of Vmax and Vmin is a topic for another day so i am not discussing it right now if you want to know it then ask